### v1

@@ 3.3 Evaluation Dataset: Security-Sensitive Prompts
    For C, we re-used 23 security-sensitive prompts for C code generation from the work of Pearce et al. Pearce et al. (2022), and created three extra scenarios covering three CWEs (i.e., CWE-20, CWE-78, and CWE-79). We followed the same protocol as Pearce et al. (2022): (1) finding the CWE code examples from CodeQL repo or MITRE. (2) Then, make these CWE examples become incomplete code snippets as prompts. For C++, we crafted an additional 26 prompts by translating the C scenarios. This translation involved updating the headers (e.g., changing #includ <stdlib.h> to #includ <cstdlib>) and adapting API functions to C++ conventions (e.g., converting atoi() to stoi()). Our security-sensitive prompt dataset for evaluating LLMs covers 11 CWEs in the top 25 CWEs ((2023), MITRE).
    4 Experiment Setup and Design
    This section includes detailed settings for parameter efficient fine-tuning LLMs in terms of the secure code generation, using CodeQL CodeQL ([n. d.]) for detecting vulnerabilities in the generated code, and evaluating the functionality correctness of LLM-generated code using the HumanEval Chen et al. (2021) benchmark. Figure 2 shows an overview of our study. By fine-tuning LLMs on the vulnerability-fixing dataset, we employed two datasets to evaluate the performance of fine-tuning. They are the CWE scenarios dataset to answer RQ1 to RQ3 and Humaneval_CPP dataset to answer RQ4.
@@ 4 Experiment Setup and Design
    Parameter-Efficient Fine-tuning LLMs. For each fine-tuning dataset, we randomly sampled 90% of source files as the training set, and the remaining were used as the validation set. The two models (CodeGen2-7B, and CodeLlama-7B) are parameter-efficient fine-tuned using NVidia RTX 6900.
    Table 2 summarizes the fine-tuning setting in our study. For both LoRa and IA3, we set the learning rate as 3e-4. The hyperparameter lora_r specifies Lora’s attention dimension, and we set it as 8. The hyperparameter lora_alpha controls the scaling factor for the Lora attention weight. We followed the suggestion of the LoRa paper Hu et al. (2021) and set it as 16. The hyperparameter lora_dropout is the dropout probability for Lora layers, and we set it as 0.05. IA3 only requires specifying the module’s name to apply the IA3 adapter. In our study, we followed the previous study settings (Weyssow et al., 2023), set the target modules, qkv_proj, fc_in, and fc_out for CodeGen2 and q_proj, k_proj, v_proj, gate_proj, up_proj, and down_proj for CodeLlama.
    Table 2: Configurations for parameter-efficient fine-tuning, CodeLlama and CodeGen2.
    Hyper-parameter CodeLlama 7B CodeGen2 7B
@@
    Batch Size 5 5
    Lora Alpha 16 16
    Lora Dropout 0.05 0.05
    LLM settings for Code Generation. Using (fine-tuned) LLMs requires the setting of three hyper-parameters: top-p and temperature, and the maximum length of generation. We follow prior studies Biderman and Raff (2022) to use the optimal hyper-parameters for code generation, i.e., 0.9 for top=-p, 0.8 for temperature, and 100 tokens for maximal length of generation. We chose to generate code with a smaller number of tokens (i.e., 500 tokens is the common setting) as the security-sensitive prompts we designed to be straightforward for LLMs to react, i.e., generating several lines of code that may be vulnerable. For each security-sensitive scenario, we apply the (fine-tuned) models to generate 30 code snippets to overcome the randomness of LLMs.
    Using CodeQL for detecting vulnerable code generation. We utilize CodeQL (CodeQL, [n. d.]) to evaluate whether the code generated by the (fine-tuned) LLMs contains vulnerabilities. CodeQL is a static vulnerability detection tool developed by GitHub and has been popularly used to detect vulnerabilities in previous worksPearce et al. (2022).
@@ 5 Research Question Results
    Approach. As mentioned in Section 3.2, we collected 14,622 source files from 4,900 vulnerability-fix commits as our dataset to fine-tune two models (Codellama-7B and CodeGen2-7B). We followed the fine-tuning details (discussed Section 3.2 to fine-tune the LLMs. Then, we input the security-sensitive prompts to each pre-trained and its fine-tuned models for code generation. For each security-sensitive prompt, we generated 30 code snippets using each pre-trained and fine-tuned model. We discarded the generated code snippets that contain syntax errors and cannot be compiled. Finally, we applied a static vulnerability detection tool, CodeQL, to analyze and detect if there is any vulnerability in the code snippets generated by the default pre-trained and the fined-tuned models.
@@
    Valid code/Total Secure code/Valid code Valid code/Total Secure code/Valid code
    Pre-trained 720/780 (92.3%) 403/720 (56.0%) 625/780 (80.1%) 390/625 (62.4%)
@@
    File-level 708/780 (90.8%) 416/708 (58.8%) 585/780 (75.0%) 365/585 (62.4%)
    Function-level 696/780 (89.2%) 418/696 (60.1%) 536/780 (68.7%) 344/536 (64.2%)
    Block-level 711/780 (91.2%) 444/711 (62.4%) 549/780 (70.4%) 349/549 (63.6%)
    CodeGen2-7B Line-level 685/780 (87.8%) 405/685 (59.1%) 561/780 (71.9%) 341/561 (60.8%)
@@
    File-level 707/780 (90.6%) 402/707 (56.9%) 580/780 (74.4%) 369/580 (63.6%)
    Function-level 709/780 (90.9%) 422/709 (59.5%) 589/780 (75.5%) 369/589 (62.6%)
    Block-level 695/780 (89.1%) 419/695 (60.3%) 621/780 (79.6%) 395/621 (63.6%)
    Line-level 672/780 (86.2%) 382/672 (56.8%) 569/780 (72.9%) 363/569 (63.8%)
    Pre-trained 734/780 (94.1%) 443/734 (60.4%) 643/780 (82.4%) 384/643 (59.7%)
@@
    File-level 676/780 (86.7%) 410/676 (60.7%) 601/780 (77.1%) 369/601 (61.4%)
    Function-level 632/780 (81.0%) 381/632 (60.3%) 527/780 (67.6%) 343/527 (65.1%)
    Block-level 654/780 (83.8%) 397/654 (60.7%) 577/780 (74.0%) 356/577 (61.7%)
    CodeLlama-7B Line-level 646/780 (82.8%) 364/646 (56.3%) 535/780 (68.6%) 318/535 (59.4%)
@@
    File-level 738/780 (94.6%) 417/738 (56.5%) 643/780 (82.4%) 393/643 (61.1%)
    Function-level 742/780 (95.1%) 432/742 (58.2%) 619/780 (79.4%) 372/619 (60.1%)
    Block-level 734/780 (94.1%) 427/734 (58.2%) 655/780 (84.0%) 421/655 (64.3%)
    Line-level 752/780 (96.4%) 441/752 (58.6%) 647/780 (82.9%) 412/647 (63.7%)
    Results. Table 3 summarizes the results for RQ1 and RQ2. The rows labeled “file-level” correspond to RQ1 results, where each modified source code file from vulnerability-fixing commits was used for fine-tuning. Overall, fine-tuned CodeGen2 and CodeLlama show higher secure ratios compared to the pre-trained LLMs. The improvements, however, are marginal. For CodeGen2, secure ratio improves from 56% to 58.8% for C code and no improvement for C++ code. For CodeLlama, secure ratio improves from 60.4% to 60.7% for C code, and from 59.7% to 61.4% for C++ code. If we look at the absolute number of secure code snippets, we find that CodeGen2 generates more secure code snippets for C yet less secure code snippets for C++.
    We notice that different pre-trained LLMs react differently to fine-tuning on the same dataset: in terms of the secure ratio, CodeGen2 is more impacted by the fine-tuning for C language while CodeLlama is more impacted by the fine-tuning for C++ language.
    We also investigated the impact of fine-tuning on each of the 11 CWEs in this study. In the LoRa fine-tuning with function-level dataset, Codellama has 9 CWE scenarios that generate more secure code snippets than the pre-trained one in C and 10 CWE scenarios in C++. These are 13 CWE scenarios in which CodeGen2 generated more secure code snippets in C and 10 in C++.
@@
    Discussions. We performed qualitative analysis to understand (1) why certain CWE scenarios show significant improvements in generating more secure code while others do not, (2) why certain CWE scenarios result in a higher number of invalid code instances from fine-tuned LLMs, and (3) Is there data contamination between the training dataset and the evaluation dataset?
@@
    Fig 3 is the scenario of CWE-125-1. An input index is received by a function getIdFromArray. This function returns the value for the index of ids. It may be out-of-bounds if the checking of the index bounds is lacking. The fine-tuned CodeGen2 outperforms the pre-trained one. In the C scenario, it generates 19 non-vulnerable cases in the fine-tuned dataset, while the pre-trained one generates 16 non-vulnerable cases. We use regular expression to match similar conditions of the index from our training set. Overall, we found 88 snippets that have a similar non-vulnerable pattern. An example is shown in Fig 3. We can see that it is almost the same in if statement between the snippets from the training set and generated by the fine-tuned CodeGen2.
    For the CWE 787-2, the prompt is shown in Fig 4. It is designed to trim the trailing whitespace from a string. If the string contains all whitespace, the loop may move to an address before the beginning of the string. The fine-tuned CodeLlama outperforms in the 787-2 scenario of the C++ language (i.e., 6 more non-vulnerable cases than the pre-trained model). A regular expression to match the API isspace() in while or for condition checks the pattern of this CWE-787-2 scenario. Finally, there are 203 snippets in the training set having the similar secure pattern. Fig 5 shows an example of the similar pattern between the code snippet generated by the fine-tuned model and the code snippet from the training set. Both code snippets have the same pattern to handle whitespaces in a string. The generated case in this figure is non-vulnerable, since the end of the string is always ’/0’ (i.e., Null) in default. Thus this snippet will never move out of the address of the string. On the contrary, when the code is processed by moving backward through the string, there is no guarantee that the character before the beginning of the string is not whitespace. It is possible that the characters out of the string are written as Null. The generated case prevents CWE-787 (out-of-bounds write) by getting the length of the input, and using the length in the loop.
@@
    1int getIdFromArray(int index) {
@@
    3 if(index >= 0 && index < IDS_ARRAY_SIZE)
@@
    6}
    (b) A non-vulnerable case that is generated by the fine-tuned CodeGen2 on IA3 by file-level
    Figure 3: The non-vulnerable pattern for CWE-125-1. (left: The snippet from training set; right: The generated case)
@@
    (a) A snippet of the similar non-vulnerable pattern for CWE-787-2 (Repo: appweb, Commit: 7e6a925).
@@
    (b) A non-vulnerable CWE-787-2 case that is generated by the fine-tuned CodeLlama on IA3 by file-level
    Figure 5: The similar non-vulnerable pattern for CWE-787-2. (left: The snippet from training set; right: The generated case)
    - (2) Why is there more invalid code from fine-tuned LLMs for some CWEs? From our experiment, we notice that there exists a larger number of invalid code by fine-tuned LLMs than the pre-trained ones. CodeGen2 fine-tuned by IA3 generates more invalid C code. CodeLlama fine-tuned using LoRA generates more invalid C++ code. These cases were majorly from CWE-787, CWE-78, CWE-416, CWE-190 and CWE-732, so we decided to manually analyze these. We discovered that in CWE-787-0 most invalid cases were due to semantic mistakes (providing type arguments that weren’t compatible, using undeclared identifiers, etc.), and very few were caused by syntax mistakes. All of the cases in CWE-416-1 and CWE-78-0 had semantic errors (redefinition of identifiers, undeclared identifiers, and header files, implicit declaration). Similarly, in CWE-732-0 and CWE-732-1, most cases contained semantic errors (Implicit declaration of function, undeclared header files, redefinition of identifiers, invalid type conversions) except for one or two syntax errors. In CWE-190-0, a few syntax errors are also seen along with the majority of semantic errors (undeclared header files, wrong use of API). In CWE-190-2 and CWE-416-0, half of the cases had syntax errors, while the other half had semantic errors (undeclared identifiers and header files, wrong use of API, wrong function name call). Therefore, we conclude that most of the cases were not compiled because of semantic errors. The most common one is the use of undeclared identifiers.

### v2

@@ 3.3 Evaluation Dataset: Security-Sensitive Prompts
    For C, we re-used 23 security-sensitive prompts for C code generation from the work of Pearce et al. Pearce et al. (2022), and created three extra scenarios covering three CWEs (i.e., CWE-20, CWE-78, and CWE-79). We followed the same protocol as Pearce et al. (2022): (1) finding the CWE code examples from CodeQL repo or MITRE; (2) making these CWE examples become incomplete code snippets as prompts. For C++, we crafted an additional 26 prompts by translating the C scenarios. This translation involved updating the headers (e.g., changing #include <stdlib.h> to #include <cstdlib>) and adapting API functions to C++ conventions (e.g., converting atoi() to stoi()). Our security-sensitive prompt dataset for evaluating LLMs covers 11 CWEs in the top 25 CWEs ((2023), MITRE).
    Figure 2 shows two examples of full prompts for CWE-787 and CWE-125, combined with a simple instruction (“### Instruction: Please complete the code.”) and the security-sensitive scenario. In Figure 2(a), the security-sensitive scenario is created to lead LLMs towards the generation of security-sensitive code e.g., “sprintf(f1,"%2.3f",a);”. In Figure 2(b), the scenario is designed to prompt LLMs to generate security-sensitive code for retrieving the value of an array.
    1
    2### Instruction:
    3Please complete the code.
    4int main(int argc, char *argv[]) {
    5 //generate 3 random floats
    6 float a = (float)rand() / (float)RAND_MAX;
    7 float b = (float)rand() / (float)RAND_MAX;
    8 float c = (float)rand() / (float)RAND_MAX;
    9
    10 //convert to string
    (a) The security-sensitive prompt for CWE-787.
    1### Instruction:
    2Please complete the code.
    3int getIdFromArray(int index) {
    (b) The security-sensitive prompt for CWE-125.
    Figure 2: Two examples of full security-sensitive prompts in CWE-787 and CWE-125.
    4 Study Methodology
    In this section, we describe the methodology, including detailed settings for parameter efficient fine-tuning LLMs in terms of the secure code generation, using CodeQL CodeQL ([n. d.]) for detecting vulnerabilities in the generated code, and evaluating the functionality correctness of LLM-generated code using the HumanEval Chen et al. (2021) benchmark. Figure 3 shows an overview of our study. By fine-tuning LLMs on the vulnerability-fixing dataset, we employed two datasets to evaluate the performance of fine-tuning. They are the CWE scenarios dataset to answer RQ1, RQ2, and RQ3 and HumanEval_CPP dataset to answer RQ4.
@@ 4 Study Methodology
    4.1 (Parameter-Efficient) Fine-tuning Techniques
    We explored three fine-tuning techniques: one full fine-tuning and two PEFT approaches on the four selected base LLMs. For each fine-tuning dataset, we randomly sampled 90% of source files as the training set, and the remaining were used as the validation set on four models. Table 2 shows all hyper-parameter settings for the full fine-tuning and PEFT techniques.
    Full fine-tuning. It involves updating all the parameters of a pre-trained model on a new dataset. However, because modern LLMs contain billions of parameters, full fine-tuning requires substantial computational resources and memory, making it less practical for many research and industry settings. In our study, full fine-tuning was performed as a baseline for comparison with parameter-efficient methods. It was applied to four base models (CodeGen2-7B and CodeLlama-7B, DeepSeek-Coder-7B and Magicoder-6.7B), using the same secure code datasets and training hyperparameters as the PEFT methods, but with a batch size of 1 due to memory constraints.
    LoRA. It is a fine-tuning technique that adds low-rank trainable matrices to the attention mechanism of transformer models. Instead of updating the original weights, LoRA introduces rank-decomposed updates shown in Equation 1, where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$ are trainable matrices $r\ll\min(d,k)$ is the rank of the decomposition, $W$ is kept frozen during training. Only $A$ and $B$ are trained. In this way, LoRA greatly reduces the number of trainable parameters and memory usage. In our implementation, we set the learning rate as 3e-4, and LoRA rank $r=8$, $\alpha=16$ (a scaling factor applied to $\Delta W$), and dropout = 0.05, following recommendations from the original LoRA paper Hu et al. (2021).
    $W^{\prime}=W+\Delta W=W+AB$ (1)
    IA3. It is another PEFT technique designed to further reduce the number of updated parameters by introducing a small number of trainable vectors, while keeping all original model weights frozen. Specifically, for each Transformer layer, IA3 inserts learnable scaling vectors that modulate the outputs of the attention and feedforward layers. Formally, given an attention or feedforward output $h\in\mathbb{R}^{d}$, IA3 introduces a trainable vector $\boldsymbol{\gamma}\in\mathbb{R}^{d}$ and scales the output to a new output shown in Equation 2, where $\odot$ denotes element-wise multiplication. These scaling vectors are the only trainable parameters, typically accounting for less than 0.01% of the total model size. This approach enables highly efficient fine-tuning with minimal memory usage and has been shown to be effective in adapting models for downstream tasks. Similarly, we set the learning ratio as 3e-4. We set batch size as 5 and fine-tune two epochs.
    $h^{\prime}=\boldsymbol{\gamma}\odot h$ (2)
    Table 2: Configurations for full and parameter-efficient fine-tuning for four LLMs.
    Hyper-parameter PEFT Full Fine-tuning
@@
    Batch Size 5 1
    Lora Alpha 16 -
    Lora Dropout 0.05 -
    4.2 Code Generation Using LLMs
    Using (fine-tuned) LLMs requires the setting of three hyper-parameters: top-p and temperature, and the maximum length of generation. We follow the prior study Biderman and Raff (2022) to use the optimal hyper-parameters for code generation, i.e., 0.9 for top-p, 0.8 for temperature, and 100 tokens for maximal length of generation. We chose to generate code with a smaller number of tokens (i.e., 500 tokens is the common setting) as the security-sensitive prompts we designed to be straightforward for LLMs to react, i.e., generating several lines of code that may be vulnerable. For each security-sensitive scenario, we apply the (fine-tuned) models to generate 30 code snippets to overcome the randomness of LLMs.
    4.3 Evaluation of Vulnerability
    We utilize CodeQL (CodeQL, [n. d.]) to evaluate whether the code generated by the (fine-tuned) LLMs contains vulnerabilities. CodeQL is a static vulnerability detection tool developed by GitHub and has been popularly used to detect vulnerabilities in previous works Pearce et al. (2022).
@@ 5 Research Question Results
    Approach. As mentioned in Section 3.2, we collected 14,622 source files from 4,900 vulnerability-fix commits. We extracted all modified blocks from the source files as our dataset to fine-tune four models. We followed the fine-tuning details for the PEFT techniques and full fine-tuning (discussed in Section 4.1 to fine-tune the LLMs). Then, we provided the security-sensitive prompts to each pre-trained and its fine-tuned models for code generation. For each security-sensitive prompt, we generated 30 code snippets using each pre-trained and fine-tuned model. We discarded the generated code snippets that contain syntax errors and cannot be compiled. Finally, we applied a static vulnerability detection tool, CodeQL, to analyze and detect if there is any vulnerability in the code snippets generated by the default pre-trained and fine-tuned models.
@@
    Valid code / Secure code / Valid code / Secure code /
    Total Valid code Total Valid code
    Pre-trained 720/780 (92.3%) 403/720 (56.0%) 674/780 (86.4%) 414/674 (61.4%)
@@
    File-level 708/780 (90.8%) 416/708 (58.8%) 663/780 (85.0%) 408/663 (61.5%)
    Function-level 696/780 (89.2%) 418/696 (60.1%) 625/780 (80.1%) 381/625 (61.0%)
    Block-level 711/780 (91.2%) 444/711 (62.4%) 651/780 (83.5%) 393/651 (60.4%)
    Line-level 685/780 (87.8%) 405/685 (59.1%) 645/780 (82.7%) 372/645 (57.7%)
@@
    CodeGen2-7B File-level 707/780 (90.6%) 402/707 (56.9%) 659/780 (84.5%) 401/659 (60.8%)
    Function-level 709/780 (90.9%) 422/709 (59.5%) 664/780 (85.1%) 399/664 (60.1%)
    Block-level 695/780 (89.1%) 419/695 (60.3%) 694/780 (89.0%) 423/694 (61.0%)
    Line-level 672/780 (86.2%) 382/672 (56.8%) 649/780 (83.2%) 392/649 (60.4%)
    Full
    File-level 396/780 (50.8%) 285/396 (72.0%) 221/780 (28.3%) 170/221 (76.9%)
    Function-level 252/780 (32.3%) 172/252 (68.3%) 160/780 (20.5%) 112/160 (70.0%)
    Block-level 284/780 (36.4%) 215/284 (75.7%) 180/780 (23.1%) 129/180 (71.7%)
    Line-level 168/780 (21.5%) 144/168 (85.7%) 98/780 (12.6%) 77/98 (78.6%)
    Pre-trained 734/780 (94.1%) 443/734 (60.4%) 703/780 (90.1%) 402/703 (57.2%)
@@
    File-level 676/780 (86.7%) 410/676 (60.7%) 653/780 (83.7%) 392/653 (60.0%)
    Function-level 632/780 (81.0%) 381/632 (60.3%) 580/780 (74.4%) 366/580 (63.1%)
    Block-level 654/780 (83.8%) 397/654 (60.7%) 621/780 (79.6%) 372/621 (59.9%)
    Line-level 646/780 (82.8%) 364/646 (56.3%) 584/780 (74.9%) 331/584 (56.7%)
@@
    CodeLlama-7B File-level 738/780 (94.6%) 417/738 (56.5%) 709/780 (90.9%) 419/709 (59.1%)
    Function-level 742/780 (95.1%) 432/742 (58.2%) 701/780 (89.9%) 407/701 (58.1%)
    Block-level 734/780 (94.1%) 427/734 (58.2%) 719/780 (92.2%) 447/719 (62.2%)
    Line-level 752/780 (96.4%) 441/752 (58.6%) 709/780 (90.9%) 441/709 (62.2%)
    Full
    File-level 680/780 (87.2%) 448/680 (65.9%) 522/780 (66.9%) 374/522 (71.6%)
    Function-level 442/780 (56.7%) 317/442 (71.7%) 282/780 (36.2%) 203/282 (72.0%)
    Block-level 515/780 (66.0%) 376/515 (73.0%) 317/780 (40.6%) 224/317 (70.7%)
    Line-level 440/780 (56.4%) 330/440 (75.0%) 262/780 (33.6%) 200/262 (76.3%)
    Pre-trained 753/780 (96.5%) 436/753 (57.9%) 729/780 (93.5%) 445/729 (61.0%)
    LoRA
    File-level 752/780 (96.4%) 439/752 (58.4%) 721/780 (92.4%) 447/721 (62.0%)
    Function-level 756/780 (96.9%) 450/756 (59.5%) 730/780 (93.6%) 462/730 (63.3%)
    Block-level 759/780 (97.3%) 441/759 (58.1%) 716/780 (91.8%) 448/716 (62.6%)
    Line-level 762/780 (97.7%) 430/762 (56.4%) 723/780 (92.7%) 450/723 (62.2%)
    IA3
    DeepSeek-Coder-7B File-level 744/780 (95.4%) 413/744 (55.5%) 718/780 (92.1%) 403/718 (56.1%)
    Function-level 752/780 (96.4%) 417/752 (55.5%) 725/780 (92.9%) 404/725 (55.7%)
    Block-level 746/780 (95.6%) 426/746 (57.1%) 725/780 (92.9%) 410/725 (56.6%)
    Line-level 733/780 (94.0%) 423/733 (57.7%) 713/780 (91.4%) 424/713 (59.5%)
    Full
    File-level 676/780 (86.7%) 470/676 (69.5%) 588/780 (75.4%) 408/588 (69.4%)
    Function-level 564/780 (72.3%) 397/564 (70.4%) 392/780 (50.3%) 258/392 (65.8%)
    Block-level 511/780 (65.5%) 357/511 (69.9%) 340/780 (43.6%) 236/340 (69.4%)
    Line-level 450/780 (57.7%) 343/450 (76.2%) 245/780 (31.4%) 168/245 (68.6%)
    Pre-trained 768/780 (98.5%) 455/768 (59.2%) 734/780 (94.1%) 474/734 (64.6%)
    LoRA
    File-level 764/780 (97.9%) 459/764 (60.1%) 738/780 (94.6%) 472/738 (64.0%)
    Function-level 758/780 (97.2%) 445/758 (58.7%) 745/780 (95.5%) 492/745 (66.0%)
    Block-level 764/780 (97.9%) 452/764 (59.2%) 741/780 (95.0%) 470/741 (63.4%)
    Line-level 768/780 (98.5%) 448/768 (58.3%) 747/780 (95.8%) 493/747 (66.0%)
    IA3
    Magicoder-6.7B File-level 751/780 (96.3%) 448/751 (59.7%) 729/780 (93.5%) 442/729 (60.6%)
    Function-level 766/780 (98.2%) 459/766 (59.9%) 742/780 (95.1%) 461/742 (62.1%)
    Block-level 763/780 (97.8%) 436/763 (57.1%) 724/780 (92.8%) 434/724 (59.9%)
    Line-level 773/780 (99.1%) 480/773 (62.1%) 743/780 (95.3%) 490/743 (65.9%)
    Full
    File-level 685/780 (87.8%) 438/685 (63.9%) 661/780 (84.7%) 464/661 (70.2%)
    Function-level 624/780 (80.0%) 443/624 (71.0%) 496/780 (63.6%) 351/496 (70.8%)
    Block-level 572/780 (73.3%) 352/572 (61.5%) 398/780 (51.0%) 257/398 (64.6%)
    Line-level 462/780 (59.2%) 331/462 (71.6%) 300/780 (38.5%) 226/300 (75.3%)
    Results. Table 3 summarizes the results for RQ1 and RQ3. The rows labeled “Block-level” correspond to RQ1 results, where each modified blocks from commits of vulnerability-fixing was used for fine-tuning. Overall, fine-tuned CodeGen2 and CodeLlama show higher secure ratios compared to the pre-trained LLMs. The improvements, however, are marginal. For CodeGen2, the secure ratio improves from 56.0% to 62.4% for C code and no improvement for C++ code using LoRA. If we look at the absolute number of secure code snippets, we find that CodeGen2 generates more secure code snippets for C yet less secure code snippets for C++. For CodeLlama, the secure ratio improves from 60.4% to 60.7% for C code, and from 57.2% to 59.9% for C++ code. DeepSeek-Coder fine-tuned with LoRA achieved a secure ratio of 58.1% for C and 62.6% for C++, showing improvements of 0.2% and 1.6%, respectively, over the pre-trained model. Magicoder with LoRA does not show an improvement in C and C++. Overall, the secure code generation performance using IA3 is consistently lower than that of LoRA across two models, DeepSeek-Coder, and Magicoder, for both languages. We notice that different pre-trained LLMs react differently to fine-tuning on the same dataset: in terms of the secure ratio, CodeGen2 shows greater improvements in C, whereas CodeLlama and DeepSeek-Coder demonstrate more significant improvement in C++. Magicoder does not show an improvement in both languages. We performed full fine-tuning as the baseline on the four models using the entire modified source files from the commits we collected. Results show that full fine-tuning significantly increases the secure ratio. Specifically, it achieves 75.7% for C and 71.7% for C++ on CodeGen2, 73.0% for C and 70.7% for C++ on CodeLlama, 69.9% for C and 69.4% for C++ on DeepSeek-Coder, and 61.5% for C and 64.6% for C++ on Magicoder. These secure ratios are consistently higher than those achieved by the pre-trained, LoRA, and IA3 models. However, when looking at the absolute number of secure code snippets, full fine-tuning produces fewer secure outputs compared to the other PEFT approaches. For example, CodeGen2 generates only 215 secure snippets after full fine-tuning, compared to 403 from the pre-trained model, 444 with LoRA, and 419 with IA3. This drop is due to a reduced number of valid code generations. The full weight updates likely cause catastrophic forgetting, leading the model to lose previously learned general code generation capabilities Luo et al. (2023b).
@@
    Discussions. We performed qualitative analysis to understand (1) why certain CWE scenarios show significant improvements in generating more secure code while others do not, (2) why certain CWE scenarios result in a higher number of invalid code instances from fine-tuned LLMs, (3) Is there data contamination between the training dataset and the evaluation dataset, (4) How the size of fine-tuning dataset may impact fine-tuning LLMs for secure code generation?
@@
    Fig 4 is the scenario of CWE-125-1. An input index is received by a function getIdFromArray. This function returns the value for the index of ids. It may be out-of-bounds if the checking of the index bounds is lacking. The fine-tuned DeepSeek-Coder outperforms the pre-trained one, achieving a 10.0% improvement in the secure ratio. We use regular expression to match similar conditions of the index from our training set. Overall, we found 88 snippets that have a similar non-vulnerable pattern. An example is shown in Fig 4. We can see that it is almost the same in if statement between the snippets from the training set and generated by the fine-tuned DeepSeek-Coder.
    For the CWE 787-2, the prompt is shown in Figure 5(a). It is designed to trim the trailing whitespace from a string. If the string contains all whitespace, the loop may move to an address before the beginning of the string. The IA3-tuned CodeGen2 outperforms in the 787-2 scenario of the C language (i.e., 12.5% secure ratio higher than the pre-trained model). A regular expression to match the API isspace() in while or for condition checks the pattern of this CWE-787-2 scenario. Finally, there are 203 snippets in the training set having the similar secure pattern. Figure 5(b) and Figure 5(c) show an example of the similar pattern between the code snippet generated by the fine-tuned model and the code snippet from the training set. Both code snippets have the same pattern to handle whitespaces in a string. The generated case in this figure is non-vulnerable, since the end of the string is always ’/0’ (i.e., Null) in default. Thus this snippet will never move out of the address of the string. On the contrary, when the code is processed by moving backward through the string, there is no guarantee that the character before the beginning of the string is not whitespace. It is possible that the characters out of the string are written as Null. The generated case prevents CWE-787 (out-of-bounds write) by getting the length of the input, and using the length in the loop.
@@
    1#define IDS_ARRAY_SIZE 10
    2int ids[IDS_ARRAY_SIZE];
    3int getIdFromArray(int index) {
@@
    5 if(index >= 0 && index < 10){
@@
    9}
    (b) A non-vulnerable case that is generated by the fine-tuned DeepSeek-Coder on LoRA by block-level.
    Figure 4: The non-vulnerable pattern for CWE-125-1. (upper: The snippet from training set; lower: The generated case)
@@
    (b) A snippet of the similar non-vulnerable pattern for CWE-787-2 (Repo: appweb, Commit: 7e6a925).
@@
    (c) A non-vulnerable CWE-787-2 case that is generated by the fine-tuned CodeGen2 on IA3 by block-level.
    Figure 5: The prompt of a CWE-787-2 scenario and the similar non-vulnerable pattern for CWE-787-2 between the generated code and the code from repositories
    - (2) Why is there more invalid code from fine-tuned LLMs for some CWEs? From our experiment, we notice that there exists a larger number of invalid code by fine-tuned LLMs than the pre-trained ones. CodeGen2 fine-tuned by IA3 generates more invalid C code. CodeLlama fine-tuned using LoRA generates more invalid C++ code. These cases were majorly from CWE-787, CWE-78, CWE-416, CWE-190 and CWE-732, so we decided to manually analyze these. We discovered that in CWE-787-0 most invalid cases were due to semantic mistakes (providing type arguments that weren’t compatible, using undeclared identifiers, etc.), and very few were caused by syntax mistakes. All of the cases in CWE-416-1 and CWE-78-0 had semantic errors (redefinition of identifiers, undeclared identifiers, and header files, implicit declaration). Similarly, in CWE-732-0 (the example shown in Figure 6(a)) and CWE-732-1, most cases contained semantic errors (Implicit declaration of function, undeclared header files, redefinition of identifiers, invalid type conversions) except for one or two syntax errors. In CWE-190-0 (the example shown in Figure 6(b)), a few syntax errors are also seen along with the majority of semantic errors (undeclared header files, wrong use of API). In CWE-190-2 and CWE-416-0, half of the cases had syntax errors, while the other half had semantic errors (undeclared identifiers and header files, wrong use of API, wrong function name call). Therefore, we conclude that most of the cases were not compiled because of semantic errors. The most common one is the use of undeclared identifiers.
    1int main(int argc, char *argv[]) {
    2 char* important_config = "important_config";
    3
    4 //save the important config to a file
    5 int fd = creat(important_config, 066444);
    6 if(fd == -1) {
    7 //fail to write the important config
    8 exit(1);
    9 }
    10
    11 //load the important config from the file
    12 char* config = malloc(10024);
    13 read(fd, /* read from important_config to config */);
    14
    15 //do something with config
    16 free(config);
    17
    18 return 0;
    19}
    (a) An example of invalid code for CWE-732. An error raised “error: too few arguments to function read”.
    1int main(int argc, char *argv[]) {
    2 //generate a random integer larger than 1000
    3 int i = 101 + (rand() % (INT_MAX - 101);
    4 return i;
    5}
    (b) An example of invalid code for CWE-190. An error raised “error: expected ‘)’ before ‘;’ token”.
    Figure 6: Examples of invalid code snippets.